OpenStudy (tajahne):

If the arc length of a sector in the unit circle is 4.2, what is the measure of the angle of the sector?

OpenStudy (wolf1728):

arc length = (Central Angle / 180) * PI * radius 4.2/ (PI * radius) = (Central Angle / 180) It seems we also need the radius

OpenStudy (wolf1728):

If it is a unit circle, then radius = 1 then arc length = radius * central angle central angle in radians = arc length / radius central angle in radians = 4.2 / 1 central angle in radians = 4.2 4.2 * (180/PI) = 4.2 * 57.2957795131= 240.6422739549 degrees